If Upper X overbar=97 sigmaσ=14and n=61​, construct a 95% confidence interval estimate of the population​ mean,
Answer #1

Here, xbar = 97 , sigma = 14 , n = 61

The z value at 95% confidence interval is,

alpha = 1 - 0.95 = 0.05
alpha/2 = 0.05/2 = 0.025
Zalpha/2 = Z0.025 = 1.96

Margin of error = E =z *(sigma/sqrt(n))
= 1.96 *(14/sqrt(61))
= 3.5133

The 95% confidence interval is

mean -E < mu < mean +E

97- 3.5133 < mu < 97+ 3.5133

93.4867 < mu < 100.5133